What is the percent composition of the element hydrogen in the compound methane, CH4?




1
Expert's answer
2022-01-28T16:19:04-0500

The molar mass of CH4 "M=12.011+4\times1.008=16.043\frac{g}{mol}"


"\%H=\frac{4\times{A_r(H)}}{M\ (CH_4)}\times100\%=\frac{4\times1.008\ g/mol}{16.043\ g/mol}\times100\%=25.13\%"


Answer: 25.13%